OpenStudy (staceyg):

what is an equation for the line with slope 2/3 and y-intercept 9?

OpenStudy (zale101):

what does the slope-intercept form look to you?

OpenStudy (staceyg):

I don't understand. Are you asking if I have a graph?

OpenStudy (staceyg):

@Zale101

OpenStudy (anonymous):

y=mx + b is your slope-intercept form m is your slope, while your y-intercept is b. Now you can easily plug in your values to get the equation :)

OpenStudy (anonymous):

y=mx+b m is 2/3 and b is 9 so y=2/3x+9.